During a turning operation on a lathe using a grooved tool, the following forces were recorded:

Group 1 2 3 4 5 6
Feed force, Fx (kN) 1.70 1.68 1.59 1.75 1.67 1.64
Radial force, Fy (kN) 1.32 1.39 1.44 1.42 1.45 1.38
Tangential force, Fz (kN) 2.18 2.37 2.38 2.32 2.29 2.33

The machine set-up includes a feed (s) of 0.24 mm/rev. and the depth of cut (d) of 2.5 mm.

Tool rake angle (\alpha) = - 6 deg.

Tool side-cutting edge angle (Cs) = 90 deg. (or Lead angle = 0 deg.)

Tool nose radius (known as corner radius) (r\varepsilon) = 0.8 mm

The work material used is AISI 1045 steel with a tensile strength, 580 MPa.

The initial workpiece diameter is 100 mm.

No cutting fluids were used during the operation.

Question:

1. Select a suitable cutting speed and the most suitable cemented carbide grade of tool material for this operation. Justify your selection

2. Calculate the metal removal rate

3. Calculate the power criterion assuming that 5.78 kW power is consumed during the operation; compare this value with that obtained using the originally selected input conditions and comment.

4. Determine the theoretical surface roughness (Ra) of the operation
Answer #1

1- for m machining of 1045 steel Ti(C,N)-Mb2C-WC-Ni best cement carbide would be best.

Resuming cutting speed v is equal to 60 metre per minute.

2- metal removal rate are equal to fdb

= .24 × 2.5 × 60 × 10^3. mm3/min

= 36 × 10^3 mm3/min .

3- cutting power ; P = Fz × V

For Fz equal to 2.18 kN

P= 2.18 kW

4- surface roughness ; h = f^2 /8R

h = (.24)^2 /8×.8 = .009 mm